arent DS × Course Hero 9 Project 5 (G)A fair survey que/X Algebra verahs.strongmind.com/courses/2724/assignments/490660?module item id-1341385 Question 1 of 4 en saved Cal Amity decided to survey 880 inhabitants of the Kingdom so that he could have a confidence level of 95% and a margin of error of 3%. He found that 72% of his sample answered yes to the second survey question, which was, Have you done anything, or will you do anything to prepare your kitchen for Gourmandos visit? Mr. Amity can infer that, most likely, between Iblank AL and _Tblank BL inhabitants of the Kingdom have done, or will do something to prepare their kitchen before Gourmandos visit Select one answer that correctly fills in blank A, and select one answer that correctly fillis in blank B, in the previous sentence blank A: 3450 U blank A:3600 9 blank A 0 blank 8: 3750 biank B:5000 blank B 3600 nB300 blank A 150 e w|曲測寵胞細鳳@gpC
rahs.strongmind.com/courses/2724/assignments/4906602module item id-1341385 esson 22 Spoiled Rotten -Lesson 23 Spoiled Rotten - Lesson 24 Spoiled Rotten-Lesson 25 Digit have been rushing around preparing food and games for Gourmandos visit, the inhabitants of the Kingdom have bee comesviolent In order to get an idea of the current attitudes of the inhabitants, and to see how they are preparing for a sit e journalist, Cal Amity, is doing a survey oximately 5000 s in the Kingdom: and other creatures, that inhabit the Kingdom, so its unreasonable for the journalist to survey all ofs of 1 to 5, with 1 beingnot terified at all; actually kind of excited in a good way and 5 being so terrified Im buying diapers one anything, or will you do anything, to prepare your kitchen for Gourmandos visic one gnything or will you do anything, to prepere any other part of your home before Gourmandos visit? Question 2 of 4 Refer to the previous questio n. consider the sample size, confidence level, and margih of errar gven in Question1 Cal Amity anything, or wil you do anything, to prepare any acher pert of your home Mr Amity can infer that most likely, between talank.A will do something to prepare another part of their homes before Gourmandos vist Encer your answers as whole numbers m the order A, B like shis 42.53 found that 24% of ns sample answered ye, tot e third survey question, which wehre you done endblank BL inhabitants of the Kingdom have done, or
Answer #1

1)
0.72 +- 0.03
= (0.69 , 0.75)

hence lower bound = 5000 * 0.69 = 3450

upper bound = 5000 * 0.75 = 3750

hence

option A) and D) are correct

Please rate

Please post next question again